OpenStudy (anonymous):

find all the zeroes and write a linear factorization of the function 5x^3-24^2+x+12

ganeshie8 (ganeshie8):

familiar wid rational root theorem ?

OpenStudy (anonymous):

no as much as I would like to be.

ganeshie8 (ganeshie8):

its okay, u heard of it during ur class right ?

OpenStudy (anonymous):

Yes. But I didn't quite understand it.

ganeshie8 (ganeshie8):

rational root theorem says this :- ``` rational roots will be of form : p/q p = factors of constant term q = factors of leading coefficient ```

ganeshie8 (ganeshie8):

here , constant term = 12 leading coefficient = 5

ganeshie8 (ganeshie8):

wat are the factors of 12 ? wat are the factors of 5 ?

OpenStudy (anonymous):

1, 2, 3, 4,5,6,12

ganeshie8 (ganeshie8):

Yes, but separate them. just list factors of 12 in one row and list factors of 5 in another row

OpenStudy (anonymous):

1,2,3,4,6,12/1,5 Like that?

ganeshie8 (ganeshie8):

yeah looks good. start testing each of them one by one

ganeshie8 (ganeshie8):

you can do that, but there is a quicker way. u knw synthetic division right ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

wonderful ! then you have all the weapons !! use it and see if we get a 0 for any of them

ganeshie8 (ganeshie8):

lets test and see if 1/1 = 1 is a zero :- 1 | 5 -24 1 12 | 5 19 20 ------------------------- 5 -19 20 42

ganeshie8 (ganeshie8):

1 is not a 0, we're not lucky yet.

OpenStudy (anonymous):

we test it on both the negative and positive factors right?

ganeshie8 (ganeshie8):

by some magic.... 4/5 looks like our guy. lets test it and see if we get a 0

ganeshie8 (ganeshie8):

yes ! more donkey work... lets finish positive stuff first, before turningto negatives

ganeshie8 (ganeshie8):

could u test and see if 4/5 gives us a 0 ? :)

OpenStudy (anonymous):

I think I did it wrong... I didn't get a 0

ganeshie8 (ganeshie8):

okay let me do this.

ganeshie8 (ganeshie8):

4/5 | 5 -24 1 12 | 4 -16 -12 ------------------------- 5 -20 -15 0 we get a 0 !! so 4/5 is one of the zero.

ganeshie8 (ganeshie8):

and the depressed equation is : 5x^2 -20x - 15

ganeshie8 (ganeshie8):

use quadratic formula and find other two zeroes

ganeshie8 (ganeshie8):

5x^2 -20x - 15 = 0 x^2 - 4x - 3 = 0

ganeshie8 (ganeshie8):

apply quadratic formula on above. (ive just cancelled 5 thru out)

OpenStudy (anonymous):

A=1 B=-4 and C=-3?

OpenStudy (anonymous):

-4plus or minus square root of 28 divided by 2 ?

ganeshie8 (ganeshie8):

Yes ! it simplifies further :o

ganeshie8 (ganeshie8):

\(\large \frac{--4\pm \sqrt{28}}{2}\) \(\large \frac{4\pm \sqrt{4 \times 7}}{2}\) \(\large \frac{4\pm 2\sqrt{7}}{2}\)

ganeshie8 (ganeshie8):

cancel 2

OpenStudy (anonymous):

4 plus or minus squareroot of 7?

ganeshie8 (ganeshie8):

\(\large \frac{4\pm 2\sqrt{7}}{2}\) \(\large \frac{2(2\pm \sqrt{7})}{2}\) \(\large 2 \pm \sqrt{7}\)

OpenStudy (anonymous):

Ah I see it now!

ganeshie8 (ganeshie8):

so the 3 zeroes are : \(\frac {4}{5}, ~~~ 2 + \sqrt{7}, ~~~~ 2 - \sqrt{7} \)

ganeshie8 (ganeshie8):

can u write the linear factorization now ?

ganeshie8 (ganeshie8):

thats just a fancy term u knw, its simple actually, we're done almost :)

ganeshie8 (ganeshie8):

linear factorization of given function is just : (x - zero)(x-zero).....

ganeshie8 (ganeshie8):

so the linear factorization of our function is :- \(\large (x-\frac {4}{5})(x-(2 + \sqrt{7}))(x-( 2 - \sqrt{7})) \)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

Okay now I understand. Thank you!

ganeshie8 (ganeshie8):

glad to hear :) yw!